# Bias in Large Language Models: Origin, Evaluation, and Mitigation

### v1

@@ 7 Conclusion
The ethical and legal implications of biased LLMs are substantial. The potential for both representational and allocational harms in real-world applications emphasizes the need for responsible AI development (Blodgett et al., 2020; Zack et al., 2024). As LLMs become increasingly integrated into diverse facets of society–from personal assistants to decision-making systems in critical sectors-—it is imperative to address these biases to prevent unintended negative consequences.
Looking forward, key research directions include developing context-sensitive bias evaluation metrics that can identify subtle biases across diverse cultural settings (Wu et al., 2024), advancing causal inference techniques for more effective debiasing methods (Li et al., 2024a), and investigating the long-term societal impacts of biased LLMs in critical domains such as healthcare and criminal justice (Obermeyer et al., 2019). Additionally, fostering interdisciplinary collaborations to address ethical challenges and develop adaptive regulatory frameworks (Gebru et al., 2021), while optimizing the balance between model performance and fairness (Joniak and Aizawa, 2022), remains crucial.
@@
Acknowledgments
A Examples of Extrinsic Biases
@@ A.1.1 Coreference Resolution
Coreference Resolution in the context of LLMs involves identifying instances where different expressions within a text refer to the same entity (Lee et al., 2017). This task is crucial for accurately interpreting the meaning of sentences, especially in cases where pronouns, names, or other referential expressions are used. The primary goal of coreference resolution is to correctly link pronouns like “he,” “she,” or “it” and definite descriptions like “the CEO” to the appropriate entity mentioned earlier in the text.
@@
    • Stereotypical Occupation Associations: In a sentence like “The doctor finished the surgery, and she went to check on the patient,” a biased model might mistakenly link “she” to a nurse or another female figure, based on the stereotype that doctors are more likely to be male. Similarly, in “The nurse finished the shift, and he went home,” the model might incorrectly associate “he” with a doctor or a male figure rather than the nurse, reflecting a bias that nurses are more likely to be female (Zhao et al., 2018a).
    • Gendered Pronoun Resolution: In a sentence like “Sam loves cooking. He is very talented in the kitchen,” a biased model might incorrectly resolve “he” to assume that Sam is male, even though the name “Sam” can be gender-neutral. This reflects a bias that cooking, when associated with talent or professionalism, is more likely to be linked to males, while in other contexts, the same task might be associated with females (Rudinger et al., 2018).
@@
    • Assumptions about Technological Proficiency: In a sentence like “The senior programmer and the young intern were debugging the code. He quickly found the bug,” a biased model might incorrectly resolve “He” to refer to the senior programmer, based on the stereotype that senior individuals are more skilled with technology, even though both the senior programmer and the young intern are equally likely candidates (Hovy and Yang, 2021).
    • Bias in Linking Pronouns to Age-related Roles: In a sentence like “The retiree and the young employee discussed the new policies. He suggested some changes,” a biased model might incorrectly resolve “he” to refer to the young employee, based on the stereotype that younger individuals are more likely to suggest changes or new ideas, whereas retirees might be stereotypically viewed as less active in professional settings (Garg et al., 2018).
@@
    • Cultural Context in Family Roles: In some cultures, family roles are strictly defined, with certain responsibilities typically assigned to specific genders or ages. A coreference resolution model might incorrectly resolve pronouns based on these cultural stereotypes. For example, in a sentence like “The eldest son took care of his siblings. She prepared dinner,” a biased model might incorrectly resolve “She” to the mother rather than the eldest daughter, based on cultural assumptions about family roles (Rudinger et al., 2018).
@@ A.1.2 Semantic Textual Similarity (STS)
    • Gendered Language and Pronoun Resolution: When comparing the similarity between sentences with gendered pronouns, such as: Sentence A: “She is a leader.”; Sentence B: “He is a leader.” A biased model might incorrectly rate the similarity between these sentences lower than it should, due to underlying gender associations, despite the fact that the sentences are nearly identical except for the pronoun (Zhao et al., 2017).
    • Gender-Stereotyped Professions: Consider two sentences: Sentence A: “The nurse administered the medication.”; Sentence B: “She took care of the patient.” If an LLM trained with gender biases in its data associates nursing primarily with women, it might assess a higher similarity between these sentences compared to another pair where the nurse is referred to as “he,” even though the gender of the nurse should not affect the similarity score (Bolukbasi et al., 2016).
@@
    • Regional Dialects and Variants: In English, the word "apartment" is commonly used in American English, while “flat” is used in British English. A model trained predominantly on one variant might incorrectly assess the similarity between “I live in an apartment” and “I live in a flat” as low, due to not recognizing the regional synonymy (Tan and Bond, 2011).
@@ A.1.3 Natural Language Inference
Natural Language Inference (NLI) is a task where the goal is to determine the relationship between a pair of sentences, typically referred to as the “premise” and the “hypothesis.” The task involves classifying the relationship between these sentences into one of three categories: entailment (the hypothesis logically follows from the premise), contradiction (the hypothesis contradicts the premise), or neutral(the hypothesis is neither supported nor contradicted by the premise; it could be true or false based on the premise alone) (Bowman et al., 2015). It is a fundamental task in understanding and reasoning with natural language and is often used to evaluate the reasoning capabilities of LLMs.
@@
    • Stereotypical Gender Roles in Professions: Given a premise like “The person is a doctor,” and a hypothesis “She is caring,” a biased model might be more likely to infer that the hypothesis is true because of the stereotype associating women with caring professions. Conversely, if the premise is “The person is an engineer,” and the hypothesis is “She is analytical,” the model might incorrectly infer this as less likely due to the stereotype that engineering is male-dominated (Rudinger et al., 2018).
    • Gendered Language and Assumptions: In a scenario where the premise is “Alex received a promotion,” and the hypothesis is “She worked very hard,” a biased model might rate this inference as less likely due to the name “Alex,” which can be gender-neutral but may be more commonly associated with males. The model may incorrectly prefer “He worked very hard” based on gender assumptions (Zhao et al., 2017).
@@
    • Bias Toward Younger Individuals in Dynamic Roles: Given the premise “The person is a dynamic entrepreneur,” and the hypothesis “The person is young,” a biased model might overly favor the hypothesis due to the stereotype that entrepreneurship and dynamism are traits associated with younger people, disregarding the possibility that older individuals can also embody these characteristics (Caliskan et al., 2017).
    • Age-Related Stereotypes in Professional Contexts: If the premise is “The CEO announced the company’s new strategic direction,” a biased model might infer that “The CEO is in their 50s” is more likely to be true (entailment), reflecting the stereotype that leadership roles are typically held by older individuals. On the other hand, it might incorrectly classify “The CEO is in their 20s” as a contradiction (Zhao et al., 2018a).
@@
    • Cultural Norms and Social Roles: Consider the premise “She is a nurse” and the hypothesis “She is caring.” A model influenced by cultural stereotypes might incorrectly infer that the hypothesis is true based on the cultural assumption that nurses are inherently caring, which reflects a bias rather than a logical inference based on the text (Rudinger et al., 2018).
    • Regional Political and Historical Contexts: A premise might state, “The Berlin Wall fell in 1989,” with a hypothesis, “Germany was reunified shortly after.” A model trained predominantly on Western narratives might correctly infer entailment, but could struggle with similar inferences in different cultural or geopolitical contexts where regional historical knowledge is less emphasized (Chen et al., 2016).
@@ A.1.4 Classification
    • Misclassification Due to Cultural Language Variants: A text classification model might incorrectly classify content written in African American Vernacular English (AAVE) as informal, unprofessional, or even toxic, due to biases against non-standard English dialects. This can result in discriminatory outcomes, such as the unfair flagging of social media posts by Black users (Sap et al., 2019a).
@@ A.1.5 Reading comprehension
    • Stereotypes in Gendered Activities: In a scenario where the passage mentions, “Alex is an excellent cook,” a biased model might assume that Alex is female and reflect this in its answers to questions about the passage. For example, when asked, "What are her skills?" the model might incorrectly use a gendered pronoun, despite "Alex" being a gender-neutral name, thereby reflecting gender stereotypes associated with cooking (Webster et al., 2018).
@@ A.2.1 Question Answering
Question Answering (QA) tasks in the context of LLMs focus on the model’s ability to provide accurate and relevant answers to questions based on a given text or knowledge base. It is an essential task in the domain of human-computer interaction, and it has been extensively used in various situations such as search engines and chatbots (Zadeh, 2006; Li et al., 2024b). The goal of QA is to extract, infer, or generate responses that directly address the query, leveraging the model’s understanding of language, context, and content (Rajpurkar et al., 2016).
@@
    • Gender Bias in Answer Generation: When asked, “What should a good leader do?”, the model might provide examples or language that align with stereotypically male attributes (e.g., “He should be assertive and decisive”), thereby implying that leadership qualities are inherently male. If asked about nurturing roles, the model might default to female pronouns and qualities, reinforcing gendered stereotypes (Zhao et al., 2018a).
@@
    • Negative Bias Toward Youth: If a question is posed like, “Are young people capable of managing a company?” a biased model might respond with, “They may lack the experience needed for such a role,” reflecting stereotypes that associate youth with inexperience, despite many young people successfully managing companies (Bolukbasi et al., 2016).
@@ A.2.2 Sentence Completions
    • Gender Bias in Descriptions of Physical Appearance: If the sentence starts with “She looked at herself in the mirror and…” a biased model might complete it with “…adjusted her makeup,” whereas “He looked at himself in the mirror and…” might be completed with “…straightened his tie.” This reflects the stereotype that women are more concerned with appearance and men with professionalism (Caliskan et al., 2017).
    • Gender Bias in Personal Attributes: If prompted with “She is very…” a model might complete the sentence with adjectives like “emotional,” “beautiful,” or “caring,” while “He is very…” might be completed with “strong,” “intelligent,” or “ambitious.” These completions reflect stereotypical views of women as being more emotional and men as being more rational or powerful (Zhao et al., 2017).
@@
    • Activity and Lifestyle Assumptions: For a prompt like “At 25, Jenny enjoys,” a biased model might complete the sentence with “partying and going out every night,” based on the stereotype that young adults are primarily interested in nightlife and socializing, ignoring the diversity of interests in this age group (Garg et al., 2018).
    • Learning and Education Stereotypes: Given the prompt “At 50, Mark decided to,” a biased model might complete it with “go back to school to finally get his degree,” assuming that older adults are “catching up” on education rather than pursuing lifelong learning or new academic challenges (Sun et al., 2019).
@@
    • Cultural Stereotyping in Sentence Completion: If the model is given a sentence like “In Japan, people often eat,” it might complete with “sushi,” reflecting a cultural stereotype that overemphasizes a specific aspect of Japanese cuisine, while ignoring the diversity of food in Japanese culture. Similarly, completing “In Mexico, people celebrate” with “Cinco de Mayo” might reinforce a limited and stereotypical understanding of Mexican culture (Hendricks et al., 2018).
    • Regional Bias in Place-Based Completions: For the sentence “In Africa, many people live in,” the model might complete with “villages,” reflecting a regional bias that assumes rural living conditions are more common across an entire continent, despite the presence of large urban areas. This completion overlooks the diversity of living environments in different African countries and regions (Birhane and Prabhu, 2021).
@@ A.2.3 Conversational
Conversational tasks in the context of LLMs involve generating, understanding, and maintaining coherent dialogue with users. These tasks are crucial for applications like chatbots, virtual assistants, and customer service automation. Conversational tasks typically require the model to not only generate contextually appropriate responses but also to track the context of the conversation, understand user intent, manage the dialogue flow, and sometimes incorporate external knowledge or personalized information (Zhang et al., 2019b).
@@
    • Stereotypical Responses Based on Gendered Prompts: When given a prompt like “Describe a nurse” versus “Describe a doctor,” an LLM might generate responses that reinforce gender stereotypes. For example, it might describe a nurse as “caring, nurturing, and female,” while describing a doctor as “authoritative, knowledgeable, and male,” despite no gender being specified in the prompts (Bolukbasi et al., 2016).
@@
    • Assumptions About Technological Savvy: In a conversation where the user is discussing technology, a biased model might assume the user is younger if they express familiarity with tech jargon or concepts. Conversely, if the user asks for help with basic technology, the model might assume the user is older and potentially provide overly simplistic explanations, which could be patronizing (Hovy and Søgaard, 2015).
    • Bias in Addressing Age-related Concerns: In a scenario where a user asks for advice on starting a new career later in life, a biased model might discourage them by emphasizing the challenges of changing careers at an older age. For instance, if the user says, “I’m thinking of switching careers at 50,” the model might respond with comments like, “It might be difficult at your age,” reflecting a bias that older individuals face more barriers in the job market (Wagner et al., 2016).
@@
    • Culturally Inappropriate Responses: A conversational LLM might provide responses that are culturally inappropriate or insensitive due to a lack of understanding of cultural norms. For example, if a user from Japan mentions “visiting a shrine,” a culturally biased model might respond with a suggestion that is more aligned with Western religious practices, failing to acknowledge the cultural significance of shrines in Japanese culture (Hovy and Spruit, 2016).

### v2

@@ 7 Conclusions
The ethical and legal implications of biased LLMs are substantial. The potential for both representational and allocational harms in real-world applications emphasizes the need for responsible AI development (Blodgett et al., 2020; Zack et al., 2024). As LLMs become increasingly integrated into diverse facets of society—from personal assistants to decision-making systems in critical sectors—it is imperative to address these biases to prevent unintended negative consequences.
Looking forward, key research directions include developing context-sensitive bias evaluation metrics that can identify subtle biases across diverse cultural settings (Wu et al., 2024), advancing causal inference techniques for more effective debiasing methods (Li et al., 2024), and investigating the long-term societal impacts of biased LLMs in critical domains such as healthcare and criminal justice (Obermeyer et al., 2019). Additionally, fostering interdisciplinary collaborations to address ethical challenges and develop adaptive regulatory frameworks (Gebru et al., 2021), while optimizing the balance between model performance and fairness (Joniak and Aizawa, 2022), remains crucial.
@@
While much of the foundational work on LLM bias has focused on earlier models, such as GPT-2, GPT-3, and BERT, a critical question is whether the observed bias trends remain, are diminished, or have evolved in more recent, state-of-the-art models such as GPT-4, Llama 3, and Claude. Preliminary evidence suggests a mixed picture. On one hand, newer models benefit from more extensive alignment procedures, including reinforcement learning from human feedback (RLHF), constitutional AI training, and red-teaming exercises, which can reduce overt toxic outputs and explicit stereotyping. For instance, Zack et al. (2024) found that GPT-4 exhibited more subtle, but still measurable, racial and gender biases in healthcare recommendations, suggesting that while the most egregious biases may be reduced, more nuanced forms persist. On the other hand, the increased scale and capability of newer models can introduce new challenges: larger context windows may amplify biases over extended interactions, and the ability to generate more fluent and persuasive text may make biased outputs harder to detect. Open-source models such as Llama 3 present additional considerations, as community fine-tuning may introduce or remove biases in ways that are difficult to track or audit. The rapid pace of model development underscores the need for continuous, longitudinal bias evaluation rather than one-time assessments, and for evaluation benchmarks that evolve alongside model capabilities.
Funding
This research received no external funding.
Data Availability
Data availability is not applicable to this review article.
Conflicts of Interest
The authors declare no conflicts of interest.
Appendix A Examples of Extrinsic Biases
@@ A.1.1 Coreference Resolution
Coreference resolution in the context of LLMs involves identifying instances where different expressions within a text refer to the same entity (Lee et al., 2017). This task is crucial for accurately interpreting the meaning of sentences, especially in cases where pronouns, names, or other referential expressions are used. The primary goal of coreference resolution is to correctly link pronouns like “he”, “she”, or “it” and definite descriptions like “the CEO” to the appropriate entity mentioned earlier in the text.
@@
    • Stereotypical occupation associations: In a sentence like “The doctor finished the surgery, and she went to check on the patient”, a biased model might mistakenly link “she” to a nurse or another female figure, based on the stereotype that doctors are more likely to be male. Similarly, in “The nurse finished the shift, and he went home”, the model might incorrectly associate “he” with a doctor or a male figure rather than the nurse, reflecting a bias that nurses are more likely to be female (Zhao et al., 2018).
    • Gendered pronoun resolution: In a sentence like “Sam loves cooking. He is very talented in the kitchen”, a biased model might incorrectly resolve “he” to assume that Sam is male, even though the name “Sam” can be gender-neutral. This reflects a bias that cooking, when associated with talent or professionalism, is more likely to be linked to males, while in other contexts, the same task might be associated with females (Rudinger et al., 2018).
@@
    • Assumptions about technological proficiency: In a sentence like “The senior programmer and the young intern were debugging the code. He quickly found the bug”, a biased model might incorrectly resolve “He” to refer to the senior programmer, based on the stereotype that senior individuals are more skilled with technology, even though both the senior programmer and the young intern are equally likely candidates (Hovy and Yang, 2021).
    • Bias in linking pronouns to age-related roles: In a sentence like “The retiree and the young employee discussed the new policies. He suggested some changes”, a biased model might incorrectly resolve “he” to refer to the young employee, based on the stereotype that younger individuals are more likely to suggest changes or new ideas, whereas retirees might be stereotypically viewed as less active in professional settings (Garg et al., 2018).
@@
    • Cultural context in family roles: In some cultures, family roles are strictly defined, with certain responsibilities typically assigned to specific genders or ages. A coreference resolution model might incorrectly resolve pronouns based on these cultural stereotypes. For example, in a sentence like “The eldest son took care of his siblings. She prepared dinner”, a biased model might incorrectly resolve “She” to the mother rather than the eldest daughter, based on cultural assumptions about family roles (Rudinger et al., 2018).
Mitigation recommendation for multilingual coreference bias: To address biases arising from pronoun omission in pro-drop languages (e.g., Chinese, Japanese, Korean, Spanish, and Arabic), we recommend training coreference resolution systems on balanced multilingual datasets that explicitly account for these grammatical peculiarities. Specifically, training data should include a substantial proportion of pro-drop language examples with gold-standard coreference annotations, so that models learn to infer referents from contextual cues rather than relying on explicit pronoun presence. Data augmentation strategies such as systematically generating pro-drop variants of existing English coreference datasets, or incorporating parallel corpora where pro-drop and non-pro-drop translations are aligned, can help models generalize across typologically diverse languages. Furthermore, evaluation benchmarks for coreference resolution should include dedicated test sets for pro-drop languages to ensure that system performance is assessed across this critical dimension of linguistic variation.
@@ A.1.2 Semantic Textual Similarity (STS)
    • Gendered language and pronoun resolution: When comparing the similarity between sentences with gendered pronouns, such as: sentence A: “She is a leader.”; sentence B: “He is a leader”. A biased model might incorrectly rate the similarity between these sentences lower than it should, due to underlying gender associations, despite the fact that the sentences are nearly identical except for the pronoun (Zhao et al., 2017).
    • Gender-stereotyped professions: Consider two sentences: sentence A: “The nurse administered the medication.”; sentence B: “She took care of the patient.” If an LLM trained with gender biases in its data associates nursing primarily with women, it might assess a higher similarity between these sentences compared to another pair where the nurse is referred to as “he”, even though the gender of the nurse should not affect the similarity score (Bolukbasi et al., 2016).
@@
    • Regional dialects and variants: In English, the word “apartment” is commonly used in American English, while “flat” is used in British English. A model trained predominantly on one variant might incorrectly assess the similarity between “I live in an apartment” and “I live in a flat” as low, due to not recognizing the regional synonymy (Tan and Bond, 2011).
@@ A.1.3 Natural Language Inference
Natural language inference (NLI) is a task where the goal is to determine the relationship between a pair of sentences, typically referred to as the “premise” and the “hypothesis”. The task involves classifying the relationship between these sentences into one of three categories: entailment (the hypothesis logically follows from the premise), contradiction (the hypothesis contradicts the premise), or neutral (the hypothesis is neither supported nor contradicted by the premise; it could be true or false based on the premise alone) (Bowman et al., 2015). This is a fundamental task in understanding and reasoning with natural language and is often used to evaluate the reasoning capabilities of LLMs.
@@
    • Stereotypical gender roles in professions: Given a premise like “The person is a doctor”, and a hypothesis “She is caring”, a biased model might be more likely to infer that the hypothesis is true because of the stereotype associating women with caring professions. Conversely, if the premise is “The person is an engineer”, and the hypothesis is “She is analytical”, the model might incorrectly infer this as less likely due to the stereotype that engineering is male-dominated (Rudinger et al., 2018).
    • Gendered language and assumptions: In a scenario where the premise is “Alex received a promotion”, and the hypothesis is “She worked very hard”, a biased model might rate this inference as less likely due to the name “Alex”, which can be gender-neutral but may be more commonly associated with males. The model may incorrectly prefer “He worked very hard” based on gender assumptions (Zhao et al., 2017).
@@
    • Bias toward younger individuals in dynamic roles: Given the premise “The person is a dynamic entrepreneur”, and the hypothesis “The person is young”, a biased model might overly favor the hypothesis due to the stereotype that entrepreneurship and dynamism are traits associated with younger people, disregarding the possibility that older individuals can also embody these characteristics (Caliskan et al., 2017).
    • Age-related stereotypes in professional contexts: If the premise is “The CEO announced the company’s new strategic direction”, a biased model might infer that “The CEO is in their 50s” is more likely to be true (entailment), reflecting the stereotype that leadership roles are typically held by older individuals. On the other hand, it might incorrectly classify “The CEO is in their 20s” as a contradiction (Zhao et al., 2018).
@@
    • Cultural norms and social roles: Consider the premise “She is a nurse” and the hypothesis “She is caring”. A model influenced by cultural stereotypes might incorrectly infer that the hypothesis is true based on the cultural assumption that nurses are inherently caring, which reflects a bias rather than a logical inference based on the text (Rudinger et al., 2018).
    • Regional political and historical contexts: A premise might state, “The Berlin Wall fell in 1989”, with a hypothesis, “Germany was reunified shortly after”. A model trained predominantly on Western narratives might correctly infer entailment, but could struggle with similar inferences in different cultural or geopolitical contexts where regional historical knowledge is less emphasized (Chen et al., 2016).
@@ A.1.4 Classification
    • Misclassification due to cultural language variants: A text classification model might incorrectly classify content written in African American Vernacular English (AAVE) as informal, unprofessional, or even toxic, due to biases against non-standard English dialects. This can result in discriminatory outcomes, such as the unfair flagging of social media posts by Black users (Sap et al., 2019).
@@ A.1.5 Reading Comprehension
    • Stereotypes in gendered activities: In a scenario where the passage mentions, “Alex is an excellent cook”, a biased model might assume that Alex is female and reflect this in its answers to questions about the passage. For example, when asked, “What are her skills?” the model might incorrectly use a gendered pronoun, despite “Alex” being a gender-neutral name, thereby reflecting gender stereotypes associated with cooking (Webster et al., 2018).
@@ A.2.1 Question Answering
Question answering (QA) tasks in the context of LLMs focus on the model’s ability to provide accurate and relevant answers to questions based on a given text or knowledge base. QA is an essential task in the domain of human–computer interaction, and it has been extensively used in various situations such as search engines and chatbots (Zadeh, 2006; Li et al., 2024). The goal of QA is to extract, infer, or generate responses that directly address the query, leveraging the model’s understanding of language, context, and content (Rajpurkar et al., 2016).
@@
    • Gender bias in answer generation: When asked, “What should a good leader do?”, the model might provide examples or language that align with stereotypically male attributes (e.g., “He should be assertive and decisive”), thereby implying that leadership qualities are inherently male. If asked about nurturing roles, the model might default to female pronouns and qualities, reinforcing gendered stereotypes (Zhao et al., 2018).
@@
    • Negative bias toward youth: If a question is posed like, “Are young people capable of managing a company?” a biased model might respond with, “They may lack the experience needed for such a role”, reflecting stereotypes that associate youth with inexperience, despite many young people successfully managing companies (Bolukbasi et al., 2016).
@@ A.2.2 Sentence Completions
    • Gender bias in descriptions of physical appearance: If the sentence starts with “She looked at herself in the mirror and…” a biased model might complete it with “…adjusted her makeup”, whereas “He looked at himself in the mirror and…” might be completed with “…straightened his tie”. This reflects the stereotype that women are more concerned with appearance and men with professionalism (Caliskan et al., 2017).
    • Gender bias in personal attributes: If prompted with “She is very…” a model might complete the sentence with adjectives like “emotional”, “beautiful”, or “caring”, while “He is very…” might be completed with “strong”, “intelligent”, or “ambitious”. These completions reflect stereotypical views of women as being more emotional and men as being more rational or powerful (Zhao et al., 2017).
@@
    • Activity and lifestyle assumptions: For a prompt like “At 25, Jenny enjoys”, a biased model might complete the sentence with “partying and going out every night”, based on the stereotype that young adults are primarily interested in nightlife and socializing, ignoring the diversity of interests in this age group (Garg et al., 2018).
    • Learning and education stereotypes: Given the prompt “At 50, Mark decided to”, a biased model might complete it with “go back to school to finally get his degree”, assuming that older adults are “catching up” on education rather than pursuing lifelong learning or new academic challenges (Sun et al., 2019).
@@
    • Cultural stereotyping in sentence completion: If the model is given a sentence like “In Japan, people often eat”, it might complete with “sushi”, reflecting a cultural stereotype that overemphasizes a specific aspect of Japanese cuisine, while ignoring the diversity of food in Japanese culture. Similarly, completing “In Mexico, people celebrate” with “Cinco de Mayo” might reinforce a limited and stereotypical understanding of Mexican culture (Hendricks et al., 2018).
    • Regional bias in place-based completions: For the sentence “In Africa, many people live in”, the model might complete with “villages”, reflecting a regional bias that assumes rural living conditions are more common across an entire continent, despite the presence of large urban areas. This completion overlooks the diversity of living environments in different African countries and regions (Birhane and Prabhu, 2021).
@@ A.2.3 Conversational
Conversational tasks in the context of LLMs involve generating, understanding, and maintaining coherent dialogue with users. These tasks are crucial for applications like chatbots, virtual assistants, and customer service automation. Conversational tasks typically require the model to not only generate contextually appropriate responses but also to track the context of the conversation, understand user intent, manage the dialogue flow, and sometimes incorporate external knowledge or personalized information (Zhang et al., 2019).
@@
    • Stereotypical responses based on gendered prompts: When given a prompt like “Describe a nurse” versus “Describe a doctor”, an LLM might generate responses that reinforce gender stereotypes. For example, it might describe a nurse as “caring, nurturing, and female”, while describing a doctor as “authoritative, knowledgeable, and male”, despite no gender being specified in the prompts (Bolukbasi et al., 2016).
@@
    • Assumptions about being tech-savvy: In a conversation where the user is discussing technology, a biased model might assume the user is younger if they express familiarity with tech jargon or concepts. Conversely, if the user asks for help with basic technology, the model might assume the user is older and potentially provide overly simplistic explanations, which could be patronizing (Hovy and Søgaard, 2015).
    • Bias in addressing age-related concerns: In a scenario where a user asks for advice on starting a new career later in life, a biased model might discourage them by emphasizing the challenges of changing careers at an older age. For instance, if the user says, “I’m thinking of switching careers at 50”, the model might respond with comments like, “It might be difficult at your age”, reflecting a bias that older individuals face more barriers in the job market (Wagner et al., 2016).
@@
    • Culturally inappropriate responses: A conversational LLM might provide responses that are culturally inappropriate or insensitive due to a lack of understanding of cultural norms. For example, if a user from Japan mentions “visiting a shrine”, a culturally biased model might respond with a suggestion that is more aligned with Western religious practices, failing to acknowledge the cultural significance of shrines in Japanese culture (Hovy and Spruit, 2016).